Q: Common stock value long Constant growth Personal Finance Problem Assume that today is January? 1,...

Q:
Common stock value long Constant growth Personal Finance Problem Assume that today is January? 1, 2013.?
?Elk County Telephone has paid the dividends shown in the following table over the previous 6? years: The? firm's dividend per share in 2013 is expected to be ?$9.
A. If you can earn 15?% on? similar-risk investments, what is the most you would be willing to pay per? share?
b.If you can earn only 12?% on? similar-risk investments, what is the most you would be willing to pay per? share?
c.Compare your findings in parts a and ?b, what is the impact of changing risk on share? value?

2012 $8.42
2011 $7.86
2010 $7.35
2009 $6.87
2008 $6.42
2007 $6.00
Answer #1

The price of a share = Expected future dividend/ required rate of return. The past dividends do not affect the price of the share.

a. Price of stock (Jan, 2013) = 9/0.15 =$60

b. Price of stock (Jan, 2013) = 9/0.12 =$75

c. As the risk decreases, so thus the required rate of return on equity, driving the price of the shares upwards.